OpenStudy (iwik):

Help plz?!?! Given the following functions f(x) and g(x), solve (f ⋅ g)(3) and select the correct answer below: f(x) = 4x^2 + 12 g(x) = x − 1 A) 47 B)−47 C)96 D)−96

OpenStudy (iwik):

@stefancipe can u help plz

OpenStudy (anonymous):

i'm sorry, but this is the first time seeing f*g, or is that f○g?

OpenStudy (iwik):

its f*g

OpenStudy (anonymous):

maybe change x in both f and g with 3 and then multiply them?

OpenStudy (anonymous):

i'll try to help you, just wait for me to find formula for multiplication :)

OpenStudy (iwik):

ok

OpenStudy (anonymous):

ok, found the solution :) try changing x in both f and g with, 3, what do you get?

OpenStudy (iwik):

like this? f(3) = 4*3^2 + 12 g(3) = 3 − 1

OpenStudy (anonymous):

so the f = 4*9+12 = 48 and g =2, then f*g equals?

OpenStudy (iwik):

-96 so D? (sry slow connection)

OpenStudy (anonymous):

wait a sec, how did you get minus?

OpenStudy (iwik):

wait 96 so its c?

OpenStudy (iwik):

am i Right?

OpenStudy (anonymous):

yes you are right

OpenStudy (anonymous):

well done